# Spatial mixing of stellar populations in globular clusters via binary–single star scattering

### v1

@@ 5 Comparison with observations
In Figs. 7 and 8, we plot the radial distributions of P1 and P2 stars in selected models at $4\,\tau_{\mathrm{rh}}$ (each with a different number of binaries) where we found the highest values of $A^{+}$. All these simulations correspond to $r_{\mathrm{h}}(0)=0.5\,\mathrm{pc}$, which is expected because the densest clusters are also the most dynamically active, and the binary–single star scattering there is more effective. In all models, P1 and P2 mix in the central regions up to the half-mass radius (i.e. $A_{1}^{+}\approx 0$, albeit the values are slightly more negative in the models with fewer or less massive binaries). The cumulative distribution functions of the model 10k-10x30Mr0.5a50 in Fig. 7 show a higher level of mixing even up to $4$ or $10\,R_{\mathrm{h}}$, proving that a higher number of binary stars can mix the central and outer populations more than fewer binaries on the same dynamical time scale. This is further supported by the bottom panels of each set of plots in the same figure, which show that the depletion of P2 stars from the core and their migration to the outer regions is dominant in the models with multiple binaries.
@@
Furthermore, in Figs. 10 and 10, we plot the time evolution of the parameter $A^{+}$ in the 10k-* and 50k-* clusters, respectively. All models have highly negative initial conditions (P2 is centrally concentrated by definition) and then gradually evolve towards a mixed state ($A^{+}{\approx}0$). We see that the more concentrated clusters and those with more binaries mix slightly more rapidly; however, all clusters with a given number of primordial binaries reach similar $A^{+}_{1}$ values at $1{-}2\,\tau_{\mathrm{rh}}$. On the other hand, the more compact clusters show higher $A^{+}_{4}$ at any given time (compare the lines of the same colour across different panels of the figure). Similarly, the models with a lower binary count require significantly longer dynamical evolution to reach the same $A^{+}_{4}$ values as the more compact ones (compare the differently coloured lines in the same panel). [^7] We note that while in the work of Pavlík et al. (2025), we showed almost full radial mixing of the clusters with $10^{4}$ stars and ten massive primordial binaries, those initial conditions corresponded to a very idealised toy model. In the models presented here, which have more realistic velocity distributions, the mixing caused by the binary is slightly suppressed by the random radial motions of the single stars. We note that our models does not showcase a clear population inversion ($A_{1}^{+}\gg 0$) anywhere up to $4.5\,\tau_{\mathrm{rh}}$.
@@
Comparing Fig. 10 and 10, the 50k-* models show lower spread of $A^{+}$ than the less populous 10k-* models. These fluctuations and inconsistencies between realisations of the same model are caused by the specific dynamical evolution and binary–binary and binary–single star interactions in each cluster, which either lead to their escape or ionisation, temporarily reducing the binary heating engine in the core (as we discussed also in terms of the radial distribution of P2 stars). We also see that the mixing of P1 and P2 is not as effective in the more populous clusters, even in the most compact ones with 20 primordial binaries. On the other hand, the trend we described above (‘more binaries equal more mixing’) is still visible. This suggests that in a more populous cluster, either a more numerous stellar-mass BH–BH subsystem is necessary, or we need to increase the BH masses – to a degree, this is also suggested by the theoretical derivations we did in Sec. 2, specifically in Fig. 2.
@@ 6 Discussion
We acknowledge that we have only explored numerically the domain of clusters with ten to fifty thousand stars, and we need to map those results to ‘real’ GCs containing up to a few million particles. The distributions of $\alpha$ (presented in Figs. 6 and 6) show that the results on binary–single star scatterings leading to stellar escapes in our models are comparable between the models, although one has five times more stars (e.g. compare the curves for 10k-30M* and 50k-30M*). This also refers back to Fig. 2, where we see that the relationship between the binary semi-major axis and the star cluster size and mass is self-similar. Therefore, this knowledge about the distribution of $\alpha$ from our less populous models can be applied to understanding what fraction of encounters would lead to the ejection from the core, the ejection from the cluster entirely, or placing the core stars in the halo in real GCs.
@@
More to the point, while the binary heating engine in our models caused the outward migration of the central P2 stars, it also pushed all the sinking P1 stars to higher radii as the cluster relaxed. We introduced three evolutionary time scales in Sec. 2 – those are 1) the time scale for individual encounters between a given binary and a random single star, 2) the time scale to eject a significant amount of P2 through binary–single star encounters from the core, and 3) the time scale for the infall of P1, which is related to the relaxation time scale of the system. To achieve the radial inversion of P1 and P2, the ejection of P2 from the core must be more rapid than the infall of P1, and once most of the P2 stars are gone from the core, it has to stop before P1 builds up in the core. If, however, the binary engine continues to scatter, or the infall of P1 happens on a shorter time scale than the ejection of the core stars, the only possible outcome is P1/P2 mixing. Although we saw in our models that any high initial binary count has been reduced over time it was never fully stopped. The kicking of stellar-mass BHs out is rather stochastic and cannot be timed or predicted from the initial conditions. Consequently, all our models mixed P1 and P2 in the long term.
An alternative mechanism to quench the binary–single star scatterings is to have a single binary system composed of two intermediate-mass black holes (IMBHs). It would be able to cause ejection of P2, but unlike stellar-mass BHs, it could also rapidly merge due to gravitational wave emission and leave the GC core due to the recoil kick (Campanelli et al., 2007; González et al., 2007) that can have up to thousands of km/s (e.g. Mahapatra et al., 2024). The scattering of P2 stars would also make the IMBH binary more eccentric, helping with merging and further amplifying the kick (Sperhake et al., 2020). For context, the general relativistic merger time of an eccentric BH–BH binary is (see, e.g. Davies et al., 2011)
@@ Acknowledgements.
This research used computational resources from e-INFRA CZ (project ID:90254), supported by the Ministry of Education, Youth and Sports of the Czech Republic; and the computational cluster Virgo at the Astronomical Institute of the Czech Academy of Sciences. V.P. is funded by the European Union’s Horizon Europe and the Central Bohemian Region under the Marie Skłodowska-Curie Actions – COFUND, Grant agreement ID:101081195 (“MERIT”). V.P. also acknowledges support from the project RVO:67985815 at the Czech Academy of Sciences. E.I.L. acknowledges support from the ERC Consolidator Grant funding scheme (project ASTEROCHRONOMETRY, https://www.asterochronometry.eu, G.A. n. 772293). A.B. acknowledges support from the Australian Research Council (ARC) Centre of Excellence for Gravitational Wave Discovery (OzGrav), through project number CE230100016. M.H. acknowledges financial support from the Excellence Cluster ORIGINS which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2094 – 390783311. A.J.W. has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101104656 and the Royal Society’s University Research Fellowship, reference URF/R1/241791.

### v2

@@ 5 Comparison with observations
In Figs. 8 and 9, we plot the radial distributions of P1 and P2 stars in selected models at $4\,\tau_{\mathrm{rh}}$ (each with a different number of binaries) where we found the highest values of $A^{+}$. All these simulations correspond to $r_{\mathrm{h}}(0)=0.5\,\mathrm{pc}$, which is expected because the densest clusters are also the most dynamically active, and the binary–single star scattering there is more effective. In all models, P1 and P2 mix in the central regions up to the half-mass radius (i.e. $A_{1}^{+}\approx 0$, albeit the values are slightly more negative in the models with fewer or less massive binaries). The cumulative distribution functions of the model 10k-10x30Mr0.5a50 in Fig. 8 show a higher level of mixing even up to $4$ or $10\,R_{\mathrm{h}}$, proving that a higher number of binary stars can mix the central and outer populations more than fewer binaries on the same dynamical time scale. This is further supported by the bottom panels of each set of plots in the same figure, which show that the depletion of P2 stars from the core and their migration to the outer regions is dominant in the models with multiple binaries.
@@
Furthermore, in Figs. 11 and 11, we plot the time evolution of the parameter $A^{+}$ in the 10k-* and 50k-* clusters, respectively. All models have highly negative initial conditions (P2 is centrally concentrated by definition) and then gradually evolve towards a mixed state ($A^{+}{\approx}0$). We see that the more concentrated clusters and those with more binaries mix slightly more rapidly; however, all clusters with a given number of primordial binaries reach similar $A^{+}_{1}$ values at $1{-}2\,\tau_{\mathrm{rh}}$. On the other hand, the more compact clusters show higher $A^{+}_{4}$ at any given time (compare the lines of the same colour across different panels of the figure). Similarly, the models with a lower binary count require significantly longer dynamical evolution to reach the same $A^{+}_{4}$ values as the more compact ones (compare the differently coloured lines in the same panel). [^7] We note that while in the work of Pavlík et al. (2025), we showed almost full radial mixing of the clusters with $10^{4}$ stars and ten massive primordial binaries, those initial conditions corresponded to a very idealised toy model. In the models presented here, which have more realistic velocity distributions, the mixing caused by the binary is slightly suppressed by the random radial motions of the single stars. We note that our models does not showcase a clear population inversion ($A_{1}^{+}\gg 0$) anywhere up to $4.5\,\tau_{\mathrm{rh}}$.
@@
Comparing Fig. 11 and 11, the 50k-* models show lower spread of $A^{+}$ than the less populous 10k-* models. These fluctuations and inconsistencies between realisations of the same model are caused by the specific dynamical evolution and binary–binary and binary–single star interactions in each cluster, which either lead to their escape or ionisation, temporarily reducing the binary heating engine in the core (as we discussed also in terms of the radial distribution of P2 stars). We also see that the mixing of P1 and P2 is not as effective in the more populous clusters, even in the most compact ones with 20 primordial binaries. On the other hand, the trend we described above (‘more binaries equal more mixing’) is still visible. This suggests that in a more populous cluster, either a more numerous stellar-mass BH–BH subsystem is necessary, or we need to increase the BH masses – to a degree, this is also suggested by the theoretical derivations we did in Sec. 2, specifically in Fig. 1.
@@ 6 Discussion
We acknowledge that we have only explored numerically the domain of clusters with ten to fifty thousand stars, and we need to map those results to ‘real’ GCs containing up to a few million particles. The distributions of $\alpha$ (presented in Figs. 7 and 7) show that the results on binary–single star scatterings leading to stellar escapes in our models are comparable between the models, although one has five times more stars (e.g. compare the curves for 10k-30M* and 50k-30M*). This also refers back to Fig. 1, where we see that the relationship between the binary semi-major axis and the star cluster size and mass is self-similar. Therefore, this knowledge about the distribution of $\alpha$ from our less populous models can be applied to understanding what fraction of encounters would lead to the ejection from the core, the ejection from the cluster entirely, or placing the core stars in the halo in real GCs.
@@
More to the point, while the binary heating engine in our models caused the outward migration of the central P2 stars, it also pushed all the sinking P1 stars to higher radii as the cluster relaxed. We introduced four evolutionary time scales in Sec. 2 – those are 1) the time scale for individual encounters between a given binary and a random single star, 2) the time scale to eject a significant amount of P2 through binary–single star encounters from the core, 3) the time scale to sufficiently mix up the stars in the core, such that all core stars have the opportunity to be ejected, which is related to the core relaxation time and 4) the time scale for the infall of P1, which is related to the relaxation time scale of the entire system. To achieve the radial inversion of P1 and P2, the ejection of P2 from the core must be more rapid than the infall of P1, and once most of the P2 stars are gone from the core, it has to stop before P1 builds up in the core. If, however, the binary engine continues to scatter, or the infall of P1 happens on a shorter time scale than the ejection of the core stars, the only possible outcome is P1/P2 mixing. Although we saw in our models that any high initial binary count has been reduced over time it was never fully stopped. The kicking of stellar-mass BHs out is rather stochastic and cannot be timed or predicted from the initial conditions. Consequently, all our models mixed P1 and P2 in the long term.
An alternative mechanism to quench the binary–single star scatterings is to have a single binary system composed of two intermediate-mass black holes (IMBHs). Let us assume one of our modelled clusters with N=50  [^0] N=50\,000 stars (where we take $0.1N$ as the core P2 stars), $r_{\mathrm{h,0}}=0.5\,\mathrm{pc}$, $m_{3}=1\,M_{\odot}$, and one BH–BH binary of $m_{1}=m_{2}=30\,M_{\odot}$. Then the ejection time scale, according to Eq. (10) and the typical value of $\alpha\approx 10^{-1.8}$ (see Fig. 7), is $\tau_{\mathrm{ej}}\approx 1.8\cdot 10^{3}\,\mathrm{Myr}$. This is very long compared to $\tau_{\mathrm{rh}}\approx 16.5\,\mathrm{Myr}$ and $\tau_{\mathrm{rc}}\approx 8\,\mathrm{Myr}$ (see the 50k-30M* panels in Figs. 3 and 3). However, when we replace the stellar-mass BH with an IMBH of $m_{1}=m_{2}=300\,M_{\odot}$, the encounter time scale from Eq. (1) becomes a thousand times smaller and $\tau_{\mathrm{ej}}\approx 1.8\,\mathrm{Myr}$. Comparing this with the 50k-10x30M* panels in Figs. 3 and 3, where the mass contained in the stellar mass BHs is equivalent to this IMBH–IMBH binary, we get $\tau_{\mathrm{rc}}\approx\tau_{\mathrm{ej}}\ll\tau_{\mathrm{rh}}$. Therefore, such IMBHs could eject the P2 stars as effectively as ten stellar-mass BH binaries. Unlike stellar-mass BHs, the IMBHs could also rapidly merge due to gravitational wave emission and leave the GC core due to the recoil kick (Campanelli et al., 2007; González et al., 2007) that can have up to thousands of km/s (e.g. Mahapatra et al., 2024). The scattering of P2 stars would also make the IMBH binary more eccentric, helping with merging and further amplifying the kick (Sperhake et al., 2020). For context, the general relativistic merger time of an eccentric BH–BH binary is (see, e.g. Davies et al., 2011)
@@ Acknowledgements.
This research used computational resources from e-INFRA CZ (project ID:90254), supported by the Ministry of Education, Youth and Sports of the Czech Republic; and the computational cluster Virgo at the Astronomical Institute of the Czech Academy of Sciences. V.P. is funded by the European Union’s Horizon Europe and the Central Bohemian Region under the Marie Skłodowska-Curie Actions – COFUND, Grant agreement ID:101081195 (“MERIT”). V.P. also acknowledges support from the project RVO:67985815 at the Czech Academy of Sciences. E.I.L. acknowledges support from the ERC Consolidator Grant funding scheme (project ASTEROCHRONOMETRY, https://www.asterochronometry.eu, G.A. n. 772293). A.B. acknowledges support from the Australian Research Council (ARC) Centre of Excellence for Gravitational Wave Discovery (OzGrav), through project number CE230100016. M.H. acknowledges financial support from the Excellence Cluster ORIGINS which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2094 – 390783311. A.J.W. has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101104656 and the Royal Society’s University Research Fellowship, reference URF/R1/241791. Last but not least, we thank the referee for valuable suggestions that improved this paper, in particular for pointing out the importance of the relative values of the core and half-mass relaxation time scales together with the time scale to eject stars from the core.